# Searching a Lightweight Network Architecture for Thermal Infrared Pedestrian Tracking

### v1

@@ III-B Implementation Details
TABLE II: Comparison of the state-of-the-art methods on the LSOTB-TIR and PTB-TIR benchmark datasets.
@@
Precision Success Norm. Precision Precision Success
DiMP [12] 0.780 0.616 0.703 0.749 0.618
@@
Ours 0.805 0.669 0.720 0.776 0.643
Figure 9: Comparison of the state-of-the-art methods on six PTB-TIR challenge subsets.
III-C Ablation Studies
III-C1 Effects of single-bottom and dual-bottom cells
The searched normal single-bottom cell is illustrated in Fig.5. To simplify architecture, only the most crucial edge is considered as input for the next layer. For instance, instead of feeding nodes S, 0, 1, and 2 into node 3, we only feed node 2 into node 3 through a 5$\times$5 dilated convolution because the architecture parameters between nodes 2 and 3 are the largest. It’s similar to attention-based pruning. The same principle applies to the reduction single-bottom cell, as shown in Fig.6. The searched normal and reduction dual-bottom cells are shown in Figs.7 and 8, respectively, which take two nodes as the input. The dual-bottom cells are more complicated than single-bottom cells. We retrain the searched architecture (random initialization) to evaluate the performance of the model. The results are shown in Table I. The single-bottom cells are very simple and have only one input, thus losing a lot of information from previous layers. Therefore, the parameters and FLOPs are small, while precision and rank are not good enough. As for the dual-bottom cells, they balance the accuracy and parameters, thus significantly improving the tracking performance.
III-C2 Effect of selecting channel randomly
We randomly select some channels and feed them into 8 operation candidates to reduce memory consumption and speed up the search process. The retrain process follows the setting of calculating all channels. We stack 16 cells (12 normal and 4 reduction cells) to form the final dual-bottom architecture. It takes 1.9 GPU days to search cells with randomly selected channels, faster than computing all channels (3.3 GPU days). The method of randomly selecting channels performs a more efficient search without compromising the performance. In this way, we can also search architecture with larger batch sizes.
III-C3 Effects of joint supervision
@@
III-D State-of-the-art Comparison
We construct a TIR-PT model that utilizes the final searched and retrained network architecture and compare it with the state-of-the-art methods, including DiMP [12], ECO-STIR [19], HSSNet [20], MCFTS [21], MDNet [22], MLSSNet [23], MMNet [5], and TADT [24], on the LSOTB-TIR [15] and PTB-TIR [16] benchmark datasets. Comparison results are shown in Table II, our method achieves competitive performance.
@@ III-D State-of-the-art Comparison
The TIR-PT model searched and retrained on the LSOTB-TIR dataset can be indeed transferable to the PTB-TIR [16] dataset, achieves the best success score of 0.641 and a comparable precision score of 0.776, outperforming DiMP with relative gains of 2.7% and 2.3%, respectively. ECO-STIR [19] uses ResNet-50 as its base network and utilizes synthetic TIR data generated from RGB data to train the model, which obtains the best precision score of 0.830, with a relative gain of 5.4% compared to our method. Still, its success score is inferior to ours, with a degradation of 2.4%. We also conduct evaluations on six challenge scenarios often occurring in pedestrian tracking, as shown in Fig.9. Our method achieves the best performance on all these challenges.
@@
IV Conclusion
In this paper, we employ a NAS approach to find an effective network architecture for the TIR-PT task. We design single-bottom and dual-bottom cells as basic searching units and randomly select channels to reduce computation and memory and feed them into operation candidates. Furthermore, the joint supervision of classification, batch hard triplet, and center loss are introduced to learn more discriminative features. The overall performance of our method outperforms the state-of-the-art methods with less computation. Extensive experiments demonstrate the effectiveness and efficiency of our proposed method.

### v2

@@ 4.3 Ablation studies
1024 0.825 0.697 0.733 23.43 M 2.67 G 3.1
8 512 0.772 0.646 0.708 1.28 M 0.39 G 0.4
1024 0.797 0.665 0.717 3.18 M 1.06 G 0.9
Table 2: Comparison of the state-of-the-art methods on the LSOTB-TIR and PTB-TIR benchmark datasets. The best two results are shown in red and blue fonts.
@@
Precision $\uparrow$ Success $\uparrow$ Norm. Precision $\uparrow$ Precision $\uparrow$ Success $\uparrow$
DBF [29] 0.770 0.625 0.703 0.839 0.626
DiMP [14] 0.739 0.621 0.668 0.737 0.589
ECO-deep [30] 0.739 0.609 0.670 0.838 0.633
@@
Ours 0.805 0.669 0.720 0.832 0.643
Figure 13: Comparison of the state-of-the-art methods on PTB-TIR challenge subsets with several attributes.
4.3.1 Effects of single-bottom and dual-bottom cells
The searched normal single-bottom cell is illustrated in Fig.8. To simplify architecture, only the most crucial edge is considered as input for the next layer. For instance, instead of feeding nodes S, 0, 1, and 2 into node 3, we only feed node 2 into node 3 through a 5$\times$5 dilated convolution because the architecture parameters between nodes 2 and 3 are the largest. It’s similar to attention-based pruning. The same principle applies to the reduction single-bottom cell, as shown in Fig.9. The searched normal and reduction dual-bottom cells are shown in Figs.10 and 11, respectively, which take two nodes as the input. The dual-bottom cells are more complicated than single-bottom cells. We retrain the searched architecture (using random initialization) to evaluate the performance of the network. The results are shown in Table 1. The single-bottom cells are very simple and have only one input, thus losing a lot of information from previous layers. Therefore, the parameters and FLOPs are small, while precision and success rates are not good enough. As for the dual-bottom cells, they balance the efficiency and parameters, thus significantly improving the tracking performance.
4.3.2 Effects of depth and width
We study the effects of architecture scaling (depth and width) by setting different configurations during the searching and training process. Results are illustrated in Table 1. Considering the efficiency, memory, and performance, we use 512 dimensions and 16 cells during searching and training. The setting allows experiments to run on a single GPU.
4.3.3 Effect of selecting channel randomly
We randomly select some channels and feed them into 8 operation candidates to reduce memory consumption and speed up the search process. The retrain process follows the setting of calculating all channels. We stack 16 cells (12 normal and 4 reduction cells) to form the final dual-bottom architecture, as shown in Fig.12. It takes 1.9 GPU days to search cells with randomly selected channels, faster than computing all channels (3.3 GPU days). The method of randomly selecting channels performs a more efficient search without compromising the performance. In this way, we can also search architecture with larger batch sizes.
4.3.4 Effects of joint supervision
@@
4.4 Comparisons with state-of-the-arts
We construct a TIR-PT model that utilizes the final searched and retrained network architecture and compare it with the state-of-the-art methods, including DBF [29], DiMP [14], ECO-stir [31], ECO-deep [30], HSSNet [19], MCFTS [18], MDNet [32], MLSSNet [20], MMNet [5], SiamFC [33], SiamTri [34], and TADT [35], on the LSOTB-TIR [25] and PTB-TIR [26] benchmark datasets. Comparison results are shown in Table 2, our method achieves competitive performance.
@@ 4.4 Comparisons with state-of-the-arts
The TIR-PT model searched and retrained on the LSOTB-TIR dataset can be indeed transferable to the PTB-TIR [26] dataset, achieves the best success score of 0.641 and a comparable precision score of 0.776, outperforming DiMP with relative gains of 2.7% and 2.3%, respectively. ECO-STIR [31] uses ResNet-50 as its base network and utilizes synthetic TIR data generated from RGB data to train the model, which obtains the best precision score of 0.830, with a relative gain of 5.4% compared to our method. Still, its success score is inferior to ours, with a degradation of 2.4%. We also conduct evaluations on challenge scenarios often occurring in pedestrian tracking, as shown in Fig.13. Our method achieves the competitive performance on all these challenges.
@@
Figure 14: Visualized comparison results of our method with the baseline DiMP tracker on six challenging video sequences from the PTB-TIR benchmark datasets. From top to bottom are classroom2, crowd3, jacket, meetion3, sidewalk3, and street3, respectively.
Figure 15: Examples of the failed cases. From left to right are circle1 and distractor1, respectively.
4.5 Visualization results
In Fig.14, we show the visualized comparison results of our method with the baseline DiMP tracker on six challenging video sequences, including classroom2, crowd3, jacket, meetion3, sidewalk3, and street3, from the PTB-TIR benchmark datasets, which utilizes ResNet-50 as its backbone network. Our method is more robust in scenes with deformation and fast motion. In comparison, DiMP is prone to be confused by background clutter and motion blur in these videos. These results indicate that the searched network architecture can learn more discriminative and robust representation. Figure 15 illustrates two instances of tracking failures. One failure results from severe occlusion by the surrounding environment, causing the backbone network to lose track of the target object and fail to recover once the occlusion ends. This suggests that our current occlusion handling mechanism is insufficient when the occlusion entirely covers the object. Severe occlusions obscure the target object’s distinctive features, complicating the tracker’s ability to maintain its trajectory. Additionally, our method struggles with objects that have similar appearances and shapes. This problem occurs because the DiMP model predictor has difficulty differentiating the target object from visually similar distractors, leading to erroneous tracking. Future work will aim to enhance occlusion and distractor handling capabilities by incorporating more advanced operational candidates.
5 Conclusion
In this paper, we employ a NAS approach to find an effective network architecture for the TIR-PT task. Our design includes single-bottom and dual-bottom cells as fundamental search units, with a strategy of random channel selection to reduce computational costs and memory usage before feeding them into operation candidates. To enhance feature discrimination, we incorporate joint supervision using classification, batch hard triplet, and center loss. Despite requiring fewer computations, our method surpasses state-of-the-art approaches in overall performance. Extensive experiments validate the effectiveness and efficiency of our proposed method. However, as our method represents an initial effort in searching for efficient network architectures for TIR-PT, the simplicity of the basic search units limits the ability to capture target-specific spatiotemporal information. Future work will focus on integrating advanced attention mechanisms from other computer vision tasks into the basic search units, aiming to learn more discriminative and informative representations for improved TIR-PT performance.